643 WA013 -10 points osColPhys2016 A ball rolls onto the path of your car as you drive down a quiet neighborhood street. To avoid hitting the chad runs to retrieve the bal, you apply your brakes for 1.40 s. The car slows down from that 15.0 m/s to 9.00 m/s. The mass of the car is 1140 kg. (a) During the time the brakes were applied, what was the average force exerted on your car? How far did the car move while braking? Reading Show My Work opeonat of 10 View Next Questovn Home My Assgnments Extension Request our last submesionis used for your score 2. O -10 points 01643 P013 1.67 mws.) The weight of an astronaut plus his space suit on the Moon is only 129 N. How much (in N) do they weigh on Earth? The acceleration due to gravity on the Moon is What is the mass (in kg) on the Moon? kg what is the mass (in kg) Show My Work 2 of 19 View Next Queston view Previous Question Question Home My Assignments Extension Roques?
Answer #1

Acceleration = change in velocity/ time = (15-9)/1.4 = 4.286 m/s2

Force F = 1140(15-9)/1.4= 4885.71 N

Force Direction is opposite to car velocity direction.

b) using v2 - u2 = 2as

2(-4.286) S = -144

S=16.8 m